1 1/4 % of __ =8.25
200% of __ =75
__ of 5.4 =2.16
__ of 1,480 =10,064
1
Expert's answer
2020-11-10T18:16:50-0500

"\implies" "\frac{11}{4}" % of __ = 8.25

"\frac{11}{4}" % of "x" = 8.25

"\frac{11}{4}"("\frac{1}{100}")("x") = 8.25

"\frac{11x}{400}" = 8.25

11"x" = 8.25(400)

11"x" = 3300

"x" = "\frac{3300}{11}" = 300

"x" = 300

"\frac{11}{4}" % of 37.5 = 8.25


"\implies" 200% 0f __ = 75

200% of "x" = 75

200("\frac{1}{100}")"x" = 75

200"x" =75(100)

200"x" = 7500

"x" = "\frac{7500}{200}" = 37.5

"x" = 37.5,

200% of 37.5 = 75

"\implies" __ of 5.4 = 2.16

"x" of 5.4 = 2.16

"x" (5.4) = 2.16

"x" = "\frac{2.16}{5.4}" = 0.4

"x" = 0.4,

0.4 of 5.4 = 2.16


"\implies" __ of 1,480 = 10,064

"x" of 1,480 = 10,064

"x" (1,480) = 10,064

"x" = "\frac{10,064}{1,480}" = 6.8

"x" = 6.8,

6.8 of 1,480 = 10,064